OpenStudy (anonymous):

I tutor a 7th grader from my school. She and I were both stumped on this problem and she needs the answer tonight. It's been so long since I've had to do this lol Convert the rate of 48 ft. per hr. to in. per min.

OpenStudy (anonymous):

48 feet per hour=48x12 inch per hour=48x12/60 inch per min=9.6 inch per min

OpenStudy (anonymous):

48 ft/ hr = 48 x 12 inches/ 60 min

OpenStudy (sleepyjess):

Oops... misread the question

OpenStudy (anonymous):

We got that far, but the teacher has asked her to put the answer in fraction form. The other problems we did like that were all set up the same way, but this one wasn't. The other answers ended up having the answer on top with a denominator of 1.

OpenStudy (anonymous):

But 9.6 is its own fraction...

OpenStudy (anonymous):

48/5

OpenStudy (sleepyjess):

576/60 = 9 36/60 = 9 3/5

OpenStudy (anonymous):

9 3/5 is the number of inches per minute?

OpenStudy (anonymous):

The feet per hour should be equivalent to the inches per minute

OpenStudy (sleepyjess):

or like @douglaswinslowcooper said 576/60 = 48/5 is the same as 9 3/5 if you want it in pure fraction form

OpenStudy (anonymous):

What units would go with 48/5 or 9 3/5?

OpenStudy (sleepyjess):

9 3/5 is 9.6 inches per minute

OpenStudy (anonymous):

Alright! Thank you!

OpenStudy (sleepyjess):

inches